OpenStudy (anonymous):

Is there any way to prove that the sum of 1/n! as n -> infinity is convergent other than to compare it with a geometric series?

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}\frac{ 1 }{ n! }\]

OpenStudy (anonymous):

The solution I have looks at the denominator first and deduces that\[n! > 2^{n-1}\] and so \[\frac{ 1 }{ n! } < \frac{ 1 }{ 2^{n-1} }\]

OpenStudy (xapproachesinfinity):

\[n!>2^{n-1}\] is not true for any n

OpenStudy (xapproachesinfinity):

there must be a condition where that holds

OpenStudy (xapproachesinfinity):

you would consider the limit of Nth term if it is zero

OpenStudy (xapproachesinfinity):

but that would turn to be difficult, i take it you are looking for a simpler way yes?

OpenStudy (anonymous):

this is a set up for the ratio test

OpenStudy (anonymous):

check that \[\lim_{n\to \infty}\frac{a_{n+1}}{a_n}<1\]

OpenStudy (anonymous):

in fact this limit is zero, since the ratio is \[\frac{n!}{(n+1)!}=\frac{1}{n+1}\]

OpenStudy (anonymous):

thank yooou =)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Another way: Suppose we know that \(e^x\) can be represented as the following series: \[e^x=\sum_{n=0}^\infty\frac{x^n}{n!}\] Setting \(x=1\) gives you a finite sum of \(e\), so the series must converge. This assumes you can establish the series converges in the first place, though.